OpenStudy (anonymous):

Suppose a parabola has a vertex (-8,-7) and also passes through the point (-7,-4). write the equation of the parabola in vertex form.

OpenStudy (klimenkov):

\(y+7=3(x+8)^2\)

OpenStudy (anonymous):

The answer is y=3(x+8) -7 .. I just dont know how to get it

OpenStudy (anonymous):

Vertex form is: \[a(x-h)^{2}+k\]

OpenStudy (anonymous):

\[(h,k) \] is your vertex

OpenStudy (anonymous):

I know that but when I plug the problem into it I don't get the correct answer. What am I doing wrong. Can you please explain how to do it to me

OpenStudy (anonymous):

So -8 and -7 are vertex

OpenStudy (anonymous):

-8 = h -7= k

OpenStudy (anonymous):

a ( -7+8) Squared + 7 is that right

OpenStudy (klimenkov):

Subtract -7 from both sides of my equation.

OpenStudy (anonymous):

No.

OpenStudy (anonymous):